How do I graph of the following:
3x+y=5
-2x+3y=4

Answer by mananth(16946)

3x+y=5................1
-2x+3y=4,,,,,,,,,,,,,,2
equation1
when x= 0, y= 5 (0,5)
when y = 0, x= 1.67 (1.67,0)
Equation 2
when x= 0 y= 1.33 (0,1.33 )
when y = 0 x= -2 (-2,0)
the solution is (1,2)